OpenStudy (anonymous):

what is the solution set to the inequality 2x^2+5x less than 12

OpenStudy (anonymous):

start with \[2x^2+5x-12<0\] then see if you can factor it

OpenStudy (anonymous):

no its wrong sorry

OpenStudy (anonymous):

ans = -4<x<3/2

OpenStudy (anonymous):

it would help maybe to actually solve it rather than write down an answer

OpenStudy (anonymous):

can you explain it to me then

OpenStudy (anonymous):

factor first you get \[(2x-3)(x+4)<0\]

OpenStudy (anonymous):

factoring i cannot explain, you just have to grind it til you find it once it is factored, you know the zeros are \(-4\) and \(\frac{3}{2}\)

OpenStudy (anonymous):

since this parabola opens up, it is negative between the zeros and positive outside them i.e. it is negative on \((-4,\frac{3}{2})\) and positive on \((-\infty, -4)\cup (\frac{3}{2},\infty)\)

OpenStudy (anonymous):

i understand the factoring

OpenStudy (anonymous):

you are asked for where it is negative so your answer is \(-4<x<\frac{3}{2}\)

OpenStudy (anonymous):

of course if you know where it is negative, you also know where it is positive as those are your only choices, unless of course it is zero

OpenStudy (anonymous):

2x^2+5x-12<0 2x^2+8x-3x-12<0 2x(x+4)-3(x+4)<0 (2x-3)(X+4)<0 -4<x<3/2 x element of (-4,3/2)

OpenStudy (anonymous):

yes if it would have asked for positive then ans would have been (−∞,−4)∪(32,∞)

OpenStudy (anonymous):

but its clearly given function <0